OpenStudy (anonymous):

Paulo is installing a circular pond in his back yard. He is surrounding the pond with decorative bricks. The pond is 4 feet in diameter. What is the length of the decorative brick border that Paulo needs to surround the pond? Use 22/7 for pi.

OpenStudy (anonymous):

C = (22/7)*5

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

@nieskaa Where do you get 5 from?

OpenStudy (anonymous):

sorry I meant 4

OpenStudy (mathstudent55):

That's better.

OpenStudy (mathstudent55):

@Nurali This is a circumference problem, not an area problem.

OpenStudy (nurali):

C = dpi C=22/7*4 C=12.5

OpenStudy (mathstudent55):

Better.